OpenStudy (anonymous):

what is the whole nunber equivalents of 20/4 ?

OpenStudy (amistre64):

20/4 = N ; multiply each side by 4 20(4)/4 = (4)N 20 = 4N ; 4 times what Number = 20?

OpenStudy (amistre64):

i see the formating aint been fixed yet

OpenStudy (amistre64):

\[\frac{20}{4}=N\text{ ; multiply each side by 4}\] \[\frac{20(4)}{4}=(4)N\text{}\] \[20=4N\text{ ; now 4 times what Number = 20?}\]

OpenStudy (amistre64):

this is all about remembering your multiplication tables :)

OpenStudy (curry):

5

OpenStudy (anonymous):

thx curry